# Electromechanical computational model of the human stomach

### v1

@@ 4.1 Influence of the smooth muscle contraction intensities
An open question in the physiology of gastric contraction waves is how the longitudinal and circumferential muscle layers coordinate to produce the large deformations observed in the human stomach during peristalsis. Competing hypotheses exist in the literature. Some studies suggest that both muscle layers contract simultaneously [60] and with similar intensity [61], while others report regionally variable or reciprocal patterns, in which the longitudinal muscle layer relaxes while the circumferential layer contracts [6]. Similar debates exist for other parts of the gastrointestinal tract [62, 63]. Overall, muscle coordination mechanisms in the gastrointestinal system remain poorly understood. For the first time, our model enables the testing of such hypotheses in the stomach through a coupled electromechanical computational framework.
@@ 4.2.1 Stomach setup
The geometry of our stomach is constructed using Coreform Cubit [59] by defining a central longitudinal path from the fundus apex to the pyloric sphincter, and define the surface geometry by ellipsoidal radii, which are shaped and scaled based on geometric comparisons of multiple magnetic resonance imaging (MRI) datasets of the human stomach. This approach allows us to capture the natural curvature and complex shape of the stomach, ensuring that the model reflects realistic anatomical features. By combining these MRI-derived dimensions with an ellipsoidal cross-section, our model preserves key anatomical characteristics, such as the curvature of the stomach, while ensuring a smoothed surface and introducing simplifications around the openings to the duodenum and esophagus. This approach ensures anatomical fidelity in these regions without introducing numerical instabilities.
The mesh resolution was guided by the convergence analysis to ensure numerical accuracy. The resulting finite element meshes comprises 210 864210\,864 linear triangular scalar transport elements for the electrophysiology problem, and 17 93917\,939 bilinear quadrilateral $7$-parameter shell elements for the solid problem. The boundary conditions applied follow those described in Section 2.5. The fully distended stomach can accommodate approximately $24\text{\,}\frac{\mathrm{l}}{}$, with intragastric pressures reaching up to $35\text{\,}\frac{\mathrm{mmHg}}{}$. Under resting conditions, intragastric pressure typically remains below $7\text{\,}\frac{\mathrm{mmHg}}{}$ [64, 65]. The stomach geometry used in our simulations provides an estimated volume of nearly $2\text{\,}\frac{\mathrm{l}}{}$. Within this physiological range, we select pressure values of $25\text{\,}\frac{\mathrm{mmHg}}{}$ in the distal stomach and $10\text{\,}\frac{\mathrm{mmHg}}{}$ in the proximal stomach to represent a partially filled state. Due to the lack of validated mixture-type material models for gastric tissue, simulation parameters are estimated to reflect physiological behavior [54] (see Table E.2). The mass fractions are assumed based on analogies with cardiac [39] and vascular tissue [66]. The stiffnesses of the spring-type boundary conditions are selected to approximate the mechanical support provided by the surrounding organs and tissues. The remaining parameters for the electrophysiology are listed in Table E.7.
@@ 4.2.2 Gastric slow wave entrainment and conduction velocity analysis
To assess whether the proposed model reproduces key features of gastric electrophysiology, we study the propagation of slow waves over a $500\text{\,}\frac{\mathrm{s}}{}$ time interval. The primary objectives are to evaluate whether the system exhibits frequency entrainment and to characterize the conduction velocity and propagation pattern of the resulting slow wave fronts. We compute the local oscillation frequencies from the transmembrane potential $v^{\textup{icc}}$ using a peak-detection algorithm applied over five successive $100\text{\,}\frac{\mathrm{s}}{}$ time windows. Only peaks with amplitudes above a threshold of $0.7$ are considered. Frequency estimation is performed for each node of the computational domain using the open-source Python library SciPy [67].
@@
Figure 9 shows the spatial distribution of these frequencies for each time window. During the initial $0100\text{\,}\frac{\mathrm{s}}{}$ window (left), the frequency field exhibits a pronounced gradient along the longitudinal direction and a smaller decrease along the circumferential direction. Over time, this gradient diminishes. By $300\text{\,}\frac{\mathrm{s}}{}$, the electrical oscillations of the ICCs synchronize to a nearly constant frequency, matching the highest intrinsic frequency of the network, approximately $3.2\text{\,}\frac{\mathrm{cpm}}{}$. This behavior indicates robust entrainment across the tissue.
@@
To further quantify the slow wave propagation, we computed the conduction velocity vector $\boldsymbol{v}_{c}$ of a slow wave front in the entrained system using a triangulation-based method [68], as detailed in C. The resulting vector fields, shown in Figure 10, confirm the aborally directed propagation originating from the pacemaker region. The pattern is anisotropic, with wave fronts spreading more rapidly circumferentially than longitudinally. Estimated velocity magnitude values range from $4\text{\,}\mathrm{mm}\text{/}\mathrm{s}$ to $6\text{\,}\mathrm{mm}\text{/}\mathrm{s}$ at the greater curvature and from $2\text{\,}\mathrm{mm}\text{/}\mathrm{s}$ to $3\text{\,}\mathrm{mm}\text{/}\mathrm{s}$ at the lesser curvature. In the pacemaker region, higher values of up to $10\text{\,}\mathrm{mm}\text{/}\mathrm{s}$ are observed. These results confirm that spatially varying diffusion coefficients yield physiologically consistent conduction velocity gradients, in agreement with experimental observations [2].
@@
To complement these velocity analyses, we generated isochrone maps that represent the activation times of the slow wave across the stomach surface, allowing the interpretation of activation patterns directly from simulation data, as comparable to techniques used in experimental gastric mapping [2]. The isochrone plots (see Figure 10) reveal multiple simultaneous slow waves, typically two to three waves propagating concurrently, reflecting the entrained, periodic activity of the gastric tissue. The spatial arrangement of the isochrones illustrates smooth and continuous physiological wavefront progression consistent with the anisotropic velocity fields, demonstrating orderly aboral propagation along the longitudinal axis toward the pylorus. The spacing between isochrones bands varies regionally, with wider spacing near the greater curvature corresponding to faster conduction, and tighter spacing near the lesser curvature indicating slower propagation. These features align well with experimental gastric electrophysiology data, further supporting the physiological relevance of our heterogeneous diffusion model.
@@ 4.2.3 Prestress
We now analyze the prestressing procedure, with a focus on determining the prestretch tensor of the ground matrix. To this end, the pressure and prestretch values of the mixture of SMCs and collagen fibers are linearly ramped up over $10$ time steps, followed by the iterative prestress algorithm to determine the ground matrix prestretch tensor (see Section 2.2.4). The algorithm terminated once the maximum Euclidean norm of the nodal displacements falls below $10^{-3}$. This tolerance value, identified by [39], offers a balance between accuracy and computational efficiency.
@@ 4.2.4 Coupled gastric motility
We now investigate the coupled electromechanical simulation of gastric motility. The first row of Figure 12 shows the spatial distribution of the normalized transmembrane potential of ICC $v^{\textup{icc}}$ at representative time points following complete entrainment of the network. At $t=$ $415\text{\,}\frac{\mathrm{s}}{}$, the pacemaker region is in its refractory phase, just before activation. Shortly thereafter, a new slow wave emerges at the pacemaker and propagates both longitudinally and circumferentially through the stomach wall. The wave fronts advances more rapidly in the circumferential direction, meeting at the lesser curvature and forming a closed activation ring. The waves directed toward the fundus dissipate within the proximal region, whereas the aborally traveling waves propagate toward the pylorus and attenuate near the sphincter. The resulting slow wave propagation pattern is periodic and stable, with two to three waves present simultaneously and an average inter-wave spacing of approximately $65\text{\,}\frac{\mathrm{mm}}{}$, consistent with experimental observations [2]. This spatial pattern reflects the successful reproduction of the native slow wave behavior in the distal stomach.
@@
Coupling this electrophysiology model to the prestressed solid mechanical model yields spatial distributions of the activation parameter $\gamma$ and displacement magnitude $\|\boldsymbol{u}\|$, shown in the second and third rows of Figure 12, respectively. The mechanical responses are tightly synchronized with the propagating slow waves, showing regionally confined contractions that follow the electrical signals of ICCs. To reproduce physiological observations, we prescribe stronger active contraction in the circumferential than in the longitudinal fiber direction ($\alpha_{\textup{c}}=0.35$ and $\alpha_{\textup{l}}=0.1$). This anisotropic activation leads to realistic, large peristaltic contractions, particularly in the distal stomach, where contraction amplitudes reach up to $8\text{\,}\frac{\mathrm{mm}}{}$, in agreement with recent in vivo measurements [13, 14].
The fourth row of Figure 12 shows temporal variations in gastric wall thickness $h_{t}$. Local increases in $h_{t}$ coincide with propagating contraction fronts, arising from the incompressibility of soft tissue and the active strain kinematics. The resulting deformation pattern is consistent with reported estimates for gastric wall deformation, although robust in vivo measurements remain limited. The overall mechanical behavior confirms the model’s ability to capture physiological gastric motility patterns on a realistic stomach geometry.
@@ 5 Conclusions
This study presents a robust and extensible computational framework for simulating coupled gastric electromechanics, capable of capturing the large, anisotropic deformations appearing during peristaltic activity on anatomically realistic geometries. A key contribution is the integration of a nonlinear, rotation-free $7$-parameter shell formulation within the open-source multiphysics code 44C. Unlike membrane elements, which lack bending stiffness and may suffer from numerical instabilities during large deformations or in regions of high curvature as common in gastric peristalsis, the proposed shell formulation ensures numerical stability while maintaining significantly lower computational cost than full 33D solid formulations. It thus offers a good balance between numerical cost, stability, and geometric fidelity for organ-scale applications.
Building on our previous electromechanical model [16], we extended the framework to include an anisotropic constrained mixture material model, prestress initialization, and non-uniform boundary conditions. These enhancements allow us to simulate physiologically realistic slow wave propagation and motility patterns, as validated by comparison to experimental data on contraction amplitudes [12, 13], conduction velocities [14], and regional frequencies [2]. Extensive convergence analyses demonstrate numerical robustness, and application to realistic stomach geometries demonstrates the framework’s suitability for whole-organ studies.
@@
A key novelty lies in the spatially heterogeneous parameterization of the model properties, such as the excitability of ICCs, the diffusion coefficient, or the fiber directions. These were assigned using smooth harmonic fields obtained from solving Laplace-Dirichlet problems aligned with anatomical axes, enabling the construction of continuous and anatomically meaningful parameter distributions. While effective, the specific parameter values remain empirically defined. This underscores the potential for future integration of data-driven methods, as pioneered in cardiac modeling [69], to improve personalization and physiological fidelity.
@@
A central insight from our parametric study is that computational modeling can elucidate physiological mechanisms that remain experimentally inaccessible. Specifically, the relative roles of circumferential and longitudinal muscle layers in gastric wall deformation are poorly characterized in vivo. Our in silico results suggest that a stronger circumferential contraction may be mechanically advantageous for effective contraction, supporting the potential physiological relevance of such configurations. This highlights the broader potential of computational models not only for prediction, but also for hypothesis generation in the absence of direct experimental data. In silico models provide a unique opportunity to investigate gastric motility processes that are difficult to assess through in vitro or in vivo approaches [54]. There is substantial inter-individual variability in stomach anatomy, including differences in shape, orientation, and spatial relation to surrounding organs and tissues. These factors result in varying mechanical boundary conditions, which may correlate with distinct contraction patterns. Such variability is difficult to capture experimentally, but can be systematically explored through computational modeling to better understand its influence on gastric biomechanics. These models can also aid in evaluating the influence of gastric geometry and contractile patterns on digestion, as well as the mechanical interplay between muscle layers and their contribution to wall deformations.
While this work represents a major step forward in gastric modeling, several limitations remain. The electrophysiology model is currently phenomenological and one-way coupled, excluding mechanosensitive feedback mechanisms known to be relevant for ICC and SMC function [70, 71]. Digesta transport is not explicitly modeled, and functional roles of anatomical regions such as the fundus, the pyloric sphincter, and the lower esophageal sphincter are not modeled in detail, despite their mechanically significance.
@@
Future extensions could focus on incorporating stretch-activated currents [6, 72], stress-assisted diffusion [73], fractional diffusivity [74], multiscale biophysical ion models [75, 76], fluid [77] and fluid–solid interaction [78]. Also contact mechanics could be investigated. Recent studies on the intestine [79] indicate that contact interactions could be important for accurately capturing gastric motility. Parameter refinement will also benefit from integration with imaging and experimental datasets through uncertainty quantification and global sensitivity analyses [80, 81].
In summary, this work provides a general, numerically robust, and physiologically informed modeling framework for gastric electromechanics on a realistic human geometry. By addressing key limitations of prior approaches, it lays the foundation for future in silico studies of the human stomach and its disorders.

### v2

@@ 4.1 Influence of the smooth muscle contraction intensities
An open question in the physiology of gastric contraction waves is how the longitudinal and circumferential muscle layers coordinate to produce the large deformations observed in the human stomach during peristalsis. Competing hypotheses exist in the literature. Some studies suggest that both muscle layers contract simultaneously [Lentle-2016-ContractileActivity] and with similar intensity [Sarna-1993-Contractions], while others report regionally variable or reciprocal patterns, in which the longitudinal muscle layer relaxes while the circumferential layer contracts [huizinga2009a]. Similar debates exist for other parts of the gastrointestinal tract [Lammers2002, Mittal2016]. Overall, muscle coordination mechanisms in the gastrointestinal system remain poorly understood. For the first time, our model enables the testing of such hypotheses in the stomach through a coupled electromechanical computational framework.
@@ 4.2.1 Stomach setup
The geometry of our stomach is constructed using Coreform Cubit [coreform2023] by defining a central longitudinal path from the fundus apex to the pyloric sphincter, and define the surface geometry by ellipsoidal radii, which are shaped and scaled based on geometric comparisons of multiple magnetic resonance imaging (MRI) datasets of the human stomach. This approach allows us to capture the natural curvature and complex shape of the stomach, ensuring that the model reflects realistic anatomical features. By combining these MRI-derived dimensions with an ellipsoidal cross-section, our model preserves key anatomical characteristics, such as the curvature of the stomach, while ensuring a smoothed surface and introducing simplifications around the openings to the duodenum and esophagus. This approach ensures anatomical fidelity in these regions without introducing numerical instabilities.
The mesh resolution was guided by the convergence analysis to ensure numerical accuracy. The resulting finite element meshes comprises 210 864210\,864 linear triangular scalar transport elements for the electrophysiology problem, and 17 93917\,939 bilinear quadrilateral $7$-parameter shell elements for the solid problem. The boundary conditions applied follow those described in Section 2.5. The fully distended stomach can accommodate approximately $24\text{\,}\frac{\mathrm{l}}{}$, with intragastric pressures reaching up to $35\text{\,}\frac{\mathrm{mmHg}}{}$. Under resting conditions, intragastric pressure typically remains below $7\text{\,}\frac{\mathrm{mmHg}}{}$ [Janssen-2011-IntragastricPressure, Wali-2013-PatientFullStomach]. The stomach geometry used in our simulations provides an estimated volume of nearly $2\text{\,}\frac{\mathrm{l}}{}$. Within this physiological range, we select pressure values of $25\text{\,}\frac{\mathrm{mmHg}}{}$ in the distal stomach and $10\text{\,}\frac{\mathrm{mmHg}}{}$ in the proximal stomach to represent a partially filled state. Due to the lack of validated mixture-type material models for gastric tissue, simulation parameters are estimated to reflect physiological behavior [Liu2024] (see Table F.2). The mass fractions are assumed based on analogies with cardiac [Gebauer-2023-CMMCardiac] and vascular tissue [Bellini-2014-MicrostructeArteialWall]. The muscle contraction intensities are set to $\alpha_{\textup{c}}=0.35$ and $\alpha_{\textup{l}}=0.1$, reflecting the relatively weaker contraction of the longitudinal muscle layer compared to the circumferential muscle layer (see Section 4.1). The stiffnesses of the spring-type boundary conditions are selected to approximate the mechanical support provided by the surrounding organs and tissues. The remaining parameters for the electrophysiology are listed in Table F.7.
@@ 4.2.2 Gastric slow wave entrainment and conduction velocity analysis
To assess whether the proposed model reproduces key features of gastric electrophysiology, we study the propagation of slow waves over a $500\text{\,}\frac{\mathrm{s}}{}$ time interval. The primary objectives are to evaluate whether the system exhibits frequency entrainment and to characterize the conduction velocity and propagation pattern of the resulting slow wave fronts. We compute the local oscillation frequencies from the transmembrane potential $v^{\textup{icc}}$ using a peak-detection algorithm applied over five successive $100\text{\,}\frac{\mathrm{s}}{}$ time windows. Only peaks with amplitudes above a threshold of $0.7$ are considered. Frequency estimation is performed for each node of the computational domain using the open-source Python library SciPy [scipy].
@@
Figure 10 shows the spatial distribution of these frequencies for each time window. During the initial $0100\text{\,}\frac{\mathrm{s}}{}$ window (left), the frequency field exhibits a pronounced gradient along the longitudinal direction and a smaller decrease along the circumferential direction. This spatial heterogeneity directly reflects the underlying excitability field $a^{\textup{icc}}$ (see Section 2.4.2), which represents regional differences in ICC intrinsic frequencies. At the beginning of the simulation, local activation patterns reflect these imposed excitability gradients, producing spatially distinct oscillation patterns. Diffusive coupling then promotes synchronization, gradually reducing the heterogeneity. By $300\text{\,}\frac{\mathrm{s}}{}$, the electrical oscillations of the ICCs synchronize to a nearly constant frequency, matching the highest intrinsic frequency of the network, approximately $3.2\text{\,}\frac{\mathrm{cpm}}{}$. This behavior indicates robust entrainment across the tissue.
@@
To further quantify the slow wave propagation, we computed the conduction velocity vector $\boldsymbol{v}_{c}$ of a slow wave front in the entrained system using a triangulation-based method [Cantwell-2015-Triangulation], as detailed in C. The resulting vector fields, shown in Figure 11, confirm the aborally directed propagation originating from the pacemaker region. The pattern is anisotropic, with wave fronts spreading more rapidly circumferentially than longitudinally. Estimated velocity magnitude values range from $4\text{\,}\mathrm{mm}\text{/}\mathrm{s}$ to $6\text{\,}\mathrm{mm}\text{/}\mathrm{s}$ at the greater curvature and from $2\text{\,}\mathrm{mm}\text{/}\mathrm{s}$ to $3\text{\,}\mathrm{mm}\text{/}\mathrm{s}$ at the lesser curvature. In the pacemaker region, higher values of up to $10\text{\,}\mathrm{mm}\text{/}\mathrm{s}$ are observed. These results confirm that spatially varying diffusion coefficients yield physiologically consistent conduction velocity gradients, in agreement with experimental observations [Grady-2021-GastricConductionReview].
@@
To complement these velocity analyses, we generated isochrone maps that represent the activation times of the slow wave across the stomach surface, allowing the interpretation of activation patterns directly from simulation data, as comparable to techniques used in experimental gastric mapping [Grady-2021-GastricConductionReview]. The isochrone plots (see Figure 11) reveal multiple simultaneous slow waves, typically two to three waves propagating concurrently, reflecting the entrained, periodic activity of the gastric tissue. The spatial arrangement of the isochrones illustrates smooth and continuous physiological wavefront progression consistent with the anisotropic velocity fields, demonstrating orderly aboral propagation along the longitudinal axis toward the pylorus. The spacing between isochrones bands varies regionally, with wider spacing near the greater curvature corresponding to faster conduction, and tighter spacing near the lesser curvature indicating slower propagation. These features align well with experimental gastric electrophysiology data, further supporting the physiological relevance of our heterogeneous diffusion model.
@@ 4.2.3 Prestress
We now analyze the prestressing procedure, with a focus on determining the prestretch tensor of the ground matrix. To this end, the pressure and prestretch values of the mixture of SMCs and collagen fibers are linearly ramped up over $10$ time steps, followed by the iterative prestress algorithm to determine the ground matrix prestretch tensor (see Section 2.2.4). The algorithm terminated once the maximum Euclidean norm of the nodal displacements falls below $10^{-3}$. This tolerance value, identified by [Gebauer-2023-CMMCardiac], offers a balance between accuracy and computational efficiency.
@@ 4.2.4 Coupled gastric motility
We now investigate the coupled electromechanical simulation of gastric motility. The first row of Figure 13 shows the spatial distribution of the normalized transmembrane potential of ICC $v^{\textup{icc}}$ at representative time points following complete entrainment of the network. At $t=$ $415\text{\,}\frac{\mathrm{s}}{}$, the pacemaker region is in its refractory phase, just before activation. Shortly thereafter, a new slow wave emerges at the pacemaker and propagates both longitudinally and circumferentially through the stomach wall. The wave fronts advances more rapidly in the circumferential direction, meeting at the lesser curvature and forming a closed activation ring. The waves directed toward the fundus dissipate within the proximal region, whereas the aborally traveling waves propagate toward the pylorus and attenuate near the sphincter. The resulting slow wave propagation pattern is periodic and stable, with two to three waves present simultaneously and an average inter-wave spacing of approximately $65\text{\,}\frac{\mathrm{mm}}{}$, consistent with experimental observations [Grady-2021-GastricConductionReview]. This spatial pattern reflects the successful reproduction of the native slow wave behavior in the distal stomach.
@@
Coupling this electrophysiology model to the prestressed solid mechanical model yields spatial distributions of the activation parameter $\gamma$ and displacement magnitude $\|\boldsymbol{u}\|$, shown in the second and third rows of Figure 13, respectively. The mechanical responses are tightly synchronized with the propagating slow waves, showing regionally confined contractions that follow the electrical signals of ICCs. To reproduce physiological observations, we prescribe stronger active contraction in the circumferential than in the longitudinal fiber direction ($\alpha_{\textup{c}}=0.35$ and $\alpha_{\textup{l}}=0.1$). This anisotropic activation leads to realistic, large peristaltic contractions, particularly in the distal stomach, where contraction amplitudes reach up to $8\text{\,}\frac{\mathrm{mm}}{}$, in agreement with recent in vivo measurements [Hosseini2023, Wang2024].
The fourth row of Figure 13 shows temporal variations in gastric wall thickness $h_{t}$. Local increases in $h_{t}$ coincide with propagating contraction fronts, arising from the incompressibility of soft tissue and the active strain kinematics. The resulting deformation pattern is consistent with reported estimates for gastric wall deformation, although robust in vivo measurements remain limited. The overall mechanical behavior confirms the model’s ability to capture physiological gastric motility patterns on a realistic stomach geometry. The coupled simulation is run on our in-house cluster ($16$ nodes with $2\times$ Intel Xeon Gold $6230$ Cascade Lake CPUs, $52$ cores, $376$ GB RAM), using $38$ cores on a single node. A representative fully coupled, entrained, and prestressed electromechanical simulation of $50\text{\,}\frac{\mathrm{s}}{}$ real time ($2500$ time steps) required approximately $2.89\text{\cdot}{10}^{05}\text{\,}\frac{\mathrm{s}}{}$ of wall-clock time ($\approx$ $80\text{\,}\frac{\mathrm{h}}{}$). Profiling data indicate that the solid mechanics solver accounts for about $76\%$ of the total runtime, while the electrophysiology solver contributes roughly $24\%$, and input/output and setup operations are negligible ($1\%$), see Table 2. This computational distribution highlights the dominant cost of the nonlinear mechanics problem and supports the adopted strategy of using a finer electrophysiology mesh coupled to a coarser mechanical discretization through the dual-mortar interface, enabling accurate yet computationally efficient electromechanical coupling.
Table 2: Profiling summary of a representative fully coupled, entrained, and prestressed electromechanical simulation of $50\text{\,}\frac{\mathrm{s}}{}$ real time (2500 time steps). The total wall-clock runtime is approximately $2.89\text{\cdot}{10}^{05}\text{\,}\frac{\mathrm{s}}{}$ ($\approx$ $80\text{\,}\frac{\mathrm{h}}{}$). Relative cost fractions illustrate the computational load distribution across solvers and operations.
Component Wall-clock time [s] Relative cost [%]
Solid mechanics solver $2.187\text{\cdot}{10}^{5}\text{\,}\frac{}{}$ $75.7$
Electrophysiology solver $7.001\text{\cdot}{10}^{4}\text{\,}\frac{}{}$ $24.2$
Input/Output and setup operations $\text{\cdot}{10}^{5}\text{\,}\frac{}{} <1<1$
Total $2.889\text{\cdot}{10}^{5}\text{\,}\frac{}{}$ $100$
@@ 5 Conclusions
Compared to our previous electromechanical model [brandstaeter2018a], the work represents a substantial extension from a conceptual prototype to a comprehensive, coupled whole-organ simulation framework for gastric electromechanics. The earlier formulation was restricted to simplified geometries, homogeneous material properties, and small, unphysiological contraction amplitudes, arising from the limitations of membrane elements. In contrast, the current framework integrates anatomically realistic geometry, advanced constitutive modeling, and spatially heterogeneous physiological parameterization to capture coordinated gastric motility with high fidelity. This study presents a robust and extensible computational framework for simulating coupled gastric electromechanics, capable of capturing the large, anisotropic deformations appearing during peristaltic activity on anatomically realistic geometries. A key contribution is the integration of a nonlinear, rotation-free $7$-parameter shell formulation within the open-source multiphysics code 44C. Unlike membrane elements, which lack bending stiffness and may suffer from numerical instabilities during large deformations or in regions of high curvature as common in gastric peristalsis, the proposed shell formulation ensures numerical stability while maintaining significantly lower computational cost than full 33D solid formulations. It thus offers a good balance between numerical cost, stability, and geometric fidelity for organ-scale applications.
Building on our previous electromechanical model [brandstaeter2018a], we extended the framework to include an anisotropic constrained mixture material model, prestress initialization, and non-uniform boundary conditions. These enhancements allow us to simulate physiologically realistic slow wave propagation and motility patterns, as validated by comparison to experimental data on contraction amplitudes [Schulze-2006-ImagingStomach, Hosseini2023], conduction velocities [Wang2024], and regional frequencies [Grady-2021-GastricConductionReview]. Specifically, the simulated contraction amplitudes ($58\text{\,}\frac{\mathrm{mm}}{}$), conduction velocities ($26\text{\,}\mathrm{mm}\text{/}\mathrm{s}$), and slow-wave frequencies ($\approx$ $3.2\text{\,}\frac{\mathrm{cpm}}{}$) fall within reported physiological ranges, supporting the biological plausibility of the model. Isochronal activation maps reproduce key experimental features, including circumferential spread near the pacemaker and longitudinal progression toward the pylorus [du2013c, OGrady-2012-AbnormalInitiationSW], demonstrating physiologically consistent slow-wave dynamics. Although gastric motility disorders are common, no direct quantitative clinical assessment of gastric wall mechanics or electrical coordination exists, as tools such as scintigraphy and electrogastrography provide only indirect measures. Physics-based computational models thus offer a means to bridge this diagnostic gap by providing mechanistic insight and quantitative predictions otherwise inaccessible.
Extensive convergence analyses demonstrate numerical robustness, and application to realistic stomach geometries demonstrates the framework’s suitability for whole-organ studies.
@@
A key novelty lies in the spatially heterogeneous parameterization of the model properties, such as the excitability of ICCs, the diffusion coefficient, or the fiber directions. These were assigned using smooth harmonic fields obtained from solving Laplace-Dirichlet problems aligned with anatomical axes, enabling the construction of continuous and anatomically meaningful parameter distributions. While effective, the specific parameter values remain empirically defined. This underscores the potential for future integration of data-driven methods, as pioneered in cardiac modeling [Barone-2020-EstimationCardiacConductivities], to improve personalization and physiological fidelity.
@@
A central insight from our parametric study is that computational modeling can elucidate physiological mechanisms that remain experimentally inaccessible. Specifically, the relative roles of circumferential and longitudinal muscle layers in gastric wall deformation are poorly characterized in vivo. Our in silico results suggest that a stronger circumferential contraction may be mechanically advantageous for effective contraction, supporting the potential physiological relevance of such configurations. This highlights the broader potential of computational models not only for prediction, but also for hypothesis generation in the absence of direct experimental data. In silico models provide a unique opportunity to investigate gastric motility processes that are difficult to assess through in vitro or in vivo approaches [Liu2024]. There is substantial inter-individual variability in stomach anatomy, including differences in shape, orientation, and spatial relation to surrounding organs and tissues. These factors result in varying mechanical boundary conditions, which may correlate with distinct contraction patterns. Such variability is difficult to capture experimentally, but can be systematically explored through computational modeling to better understand its influence on gastric biomechanics. These models can also aid in evaluating the influence of gastric geometry and contractile patterns on digestion, as well as the mechanical interplay between muscle layers and their contribution to wall deformations.
While this work represents a major step forward in gastric modeling, several limitations remain. The electrophysiology model is currently phenomenological and one-way coupled, excluding mechanosensitive feedback mechanisms known to be relevant for ICC and SMC function [Kraichely-2007-MechanosensitiveIonChannels, Huizinga-2018-StretchCurrents]. Further, fluid–structure interactions are not explicitly modeled. A spatially uniform intragastric pressure is used to approximate the mechanical load on the gastric wall, providing a first-order approximation of wall mechanics but neglecting local fluid effects, shear stresses, and feedback between fluid motion and tissue deformation. This simplification may limit the ability to capture fully physiologically accurate patterns of intraluminal flow and wall deformation, representing an important avenue for future research [Brandstaeter2019, Liu2024]. In addition, although our framework, combining the $7$-parameter shell formulation with the constrained mixture model, efficiently captures large deformations and tissue mechanics, it represents the gastric wall in a homogenized manner. Intramural heterogeneity arising from the multilayered structure and its fiber dispersion is not explicitly resolved, which limits the reproduction of intramural mechanics. Digesta transport is not explicitly modeled, and functional roles of anatomical regions such as the fundus, the pyloric sphincter, and the lower esophageal sphincter are not modeled in detail, despite their mechanically significance.
@@
Future extensions could focus on incorporating stretch-activated currents [huizinga2009a, Joshi2021], stress-assisted diffusion [Cherubini2017], fractional diffusivity [Cusimano2019], multiscale biophysical ion models [Mah2020, Athavale-2024-RatGastricSlowWave], fluid [Palmada2023] and fluid–solid interaction [Fuchs2021]. Also contact mechanics could be investigated. Recent studies on the intestine [Djoumessi2025] indicate that contact interactions could be important for accurately capturing gastric motility. Parameter refinement will also benefit from integration with imaging and experimental datasets through uncertainty quantification and global sensitivity analyses [Brandstaeter2021a, Wirthl2023].
In summary, this work provides a general, numerically robust, and physiologically informed modeling framework for gastric electromechanics on a realistic human geometry. Beyond methodological innovation, the framework provides a step toward connecting computational modeling with clinical applications by offering a quantitative tool to analyze normal and pathological gastric function. The modular design of the numerical framework allows for straightforward incorporation of more detailed electrophysiological cell models, alternative constitutive formulations, or more advanced boundary conditions, offering a flexible platform for future investigations. Furthermore, such simulations can support clinical interpretation of motility imaging, assist in virtual testing of pacing therapies, and provide mechanistic insight into disorders such as gastroparesis or functional dyspepsia. Integration with clinical imaging and electrogastrography could enable patient-specific model calibration and non-invasive assessment of gastric function, paving the way toward translational and personalized modeling of gastrointestinal motility. By addressing key limitations of prior approaches, it lays the foundation for future in silico studies of the human stomach and its disorders.